# On the choice of the non-trainable internal weights in random feature maps

### v1

@@ 5.1. Effect of the quality of internal weights on the forecast time $\tau_{f}$
In this section we investigate how the forecasting skill of a random feature surrogate model (3) improves with increasing the number of good rows $N_{g}$. We would like to have internal parameters resulting in large mean forecast times $\tau_{f}$ with relatively small standard deviations. For chaotic dynamical systems we expect a residual variance of the forecast time due to the sensitivity to small changes in the model: small changes in the internal parameters may cause the surrogate models to deviate from each other after some time.
@@
Figure 8 shows the dependence of the mean forecast time $\mathbb{E}[\tau_{f}]$ and the associated coefficient of variation $\sigma[\tau_{f}]/\mathbb{E}[\tau_{f}]$ on $p_{g}$ for various values of the feature dimensions $D_{r}$ and training data lengths $N$. It is clearly seen that increasing the number of good rows increases the mean forecast time and decreases the coefficient of variation as desired. As expected, for fixed feature dimension $D_{r}$ increasing the training data length $N$ is beneficial. Similarly, for fixed training data length $N$, increasing the feature dimension $D_{r}$ is beneficial. The observation that, for fixed data length $N$, the mean forecast time $\mathbb{E}[\tau_{f}]$ saturates upon increasing $p_{g}$ once a sufficiently large number of good features $N_{g}=p_{g}D_{r}$ are present, suggests that the distribution of the forecast time $\tau_{f}$ converges reflecting a residual uncertainty of the chaotic surrogate model. This is confirmed in Figure 9 where we see convergence of the empirical histograms of the forecast time for increasing values of $D_{r}$ in the case when $p_{g}=1$.
@@
We briefly discuss the effect of the regularization parameter $\beta$ on the forecasting skill. We show in Figure 12 the mean forecast time $\mathbb{E}[\tau_{f}]$ and coefficient of variation $\sigma[\tau_{f}]/\mathbb{E}[\tau_{f}]$ as a function of $p_{g}$ for a range of regularization parameters $\beta\in[2^{-25},2^{-13}]$. For fixed feature dimension $D_{r}=300$, we see that $\beta=2^{-19}$ is optimal within this range in terms of the mean forecast time (left panel) and the coefficient of variation (right panel) once sufficiently many good features are present with $p_{g}>0.33$. Note that we had previously employed $\beta=4\times 10^{-5}\approx 2^{-14.6}$.
@@ 5.2. Effect of the quality of internal weights on the outer weights $\mathbf{W}$
We begin by recording the Frobenius norm $\mathbf{\|W\|}$ of the learned outer weights as a function of the fraction of good features $p_{g}$ (bad features are again roughly equally distributed between linear and saturated features). Figure 13 shows the mean of $\mathbf{\|W\|}$ as a function of $p_{g}$ on a $\log$-$\log$ scale for the simulations used in Figure 8. It is seen that $\|\mathbf{W}\|$ is a decreasing function of the number of good features. The solution of the linear regression problem $\mathbf{W}$ minimizes the loss function (10). Once there are sufficiently many good features, the training data can be sufficiently well fit, decreasing the first term of the loss function. Increasing the number of good features further then allows to decrease the second regularizing term of the loss function, leading to a decrease of $\|\mathbf{W}\|$. Assuming that the true one-step map $\Psi_{\Delta t}$ in (2) lies in the domain of the random feature map (3) with infinitely many features, the first term of the loss function should scale with the usual Monte-Carlo estimate scaling of $\mathcal{O}(1/D_{r})$, suggesting a scaling of the regularization term $\|\mathbf{W}\|\sim 1/\sqrt{D_{r}}$. In the right panel of Figure 13 we show that the mean of $\|\mathbf{W}\|$ roughly scales as $\|\mathbf{W}\|\sim 1/D_{r}^{0.54}$ when all the internal weights correspond to good features with $p_{g}=1$, suggesting that the true one-step map can be well approximated by random features with a $\tanh$-activation function. We remark that the Monte-Carlo scaling is valid for $D_{r}>256$ only, i.e. provided sufficiently many good features are present.
@@
Figure 13. Left: The mean of the Frobenius norm of the outer weights, $\|\mathbf{W}\|$, as a function of $p_{g}$ on a log-log scale. Right: The mean of the Frobenius norm of the outer weights, $\|\mathbf{W}\|$, as a function of the feature dimension $D_{r}$ for $p_{g}=1$. Shown is also a line of best fit with approximate slope $-0.54$. Expectations were taken over $500$ realizations.
@@
We now investigate how the decrease in the outer weights $\mathbf{W}$ is distributed over the various features. We will see that the outer weights are learned to suppress the bad features provided there are sufficiently many good features allowing for a reduction of the loss function. Let us denote the $i$-th column of $\mathbf{W}$ by $\mathbf{W}^{i}$. The columns $\mathbf{W}^{i}$ are the weights attributed to the features produced by the $i$-th row of the internal weights $(\mathbf{w}_{i},b_{i})$. We expect the outer weights corresponding to good rows to be significantly larger than those corresponding to bad rows. We label columns of $\mathbf{W}$ that have only small entries with absolute value smaller than a threshold $\delta=1$ by $\mathcal{N}^{0}$.
@@
To study the suppression of bad features by such small columns of the learned outer weights $\mathbf{W}$, we design two sets of numerical experiments: one in which bad features are entirely comprised of linear features and one in which bad features are entirely comprised of saturated features.
@@
In the first set we initialize a random feature map with $D_{r}=300$ features consisting of only bad linear features. We then successively replace one linear feature by a good feature, i.e. replacing one inner linear weight row $(\mathbf{w}_{i}^{\rm in},b_{i}^{\rm in})$ by a good row. At each step we record the corresponding linear regression solution $\mathbf{W}$. Figure 14 shows the normalized supremum norm of columns of $\mathbf{W}$ after $N_{g}=10$, $N_{g}=50$ and $N_{g}=150$ bad linear features have been replaced by good features. The red dots signify small columns which do not contain any entry with absolute value larger than $\delta=1$. It is clearly seen that linear features are suppressed by the columns of $\mathbf{W}$. Note that not all linear features are entirely suppressed.
@@
In the second experiment we follow the same procedure as before except we start with only saturated random features. In Figure 15 it is seen that saturated features are suppressed even stronger by the outer weights than linear features. In contrast to linear features, saturated features are effectively fully suppressed once the number of good features exceeds $N_{g}=50$.
@@
Figure 14. Normalized supremum norm of the columns of $\mathbf{W}$ for different numbers of good features with $N_{g}=10$, $N_{g}=50$ and $N_{g}=150$ and otherwise exclusively linear bad features. The $x$-axis represents column indices. The good and linear columns are indicated in blue and orange, respectively. The red dots signify columns with supremum norm less than $\delta=1$. The overall feature dimension is $D_{r}=300$ and the outer weights were obtained from training data of length $N=20,000$.
@@
Figure 15. Normalized supremum norm of the columns of $\mathbf{W}$ for different numbers of good features with $N_{g}=10$, $N_{g}=50$ and $N_{g}=150$ and otherwise exclusively saturated bad features. The $x$-axis represents column indices. The good and saturated columns are indicated in blue and green, respectively. The red dots signify columns with supremum norm less than $\delta=1$. The overall feature dimension is $D_{r}=300$ and the outer weights were obtained from training data of length $N=20,000$.
@@ 6. Comparison with a single-layer feedforward network trained with gradient descent
$\mathbf{\Theta}=\underset{\mathbf{W}_{\rm in},\mathbf{b}_{\rm in},\mathbf{W}}{\operatorname*{arg\,min}}\;\mathcal{L}(\mathbf{U}),$ (32)
with $\mathbf{\Theta}=(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in},\mathbf{W})$ and the loss function $\mathcal{L}$ defined in (10). To solve the optimization problem (32) we employ gradient descent. In order to fairly compare with the results from the random feature model, we fix the width of the internal layer to $D_{r}=300$ and employ a regularization parameter of $\beta=4\times 10^{-5}$. We use training data of length $N=20,000$. To initialize the network weights we use the standard Glorot initialization [11]. We use an adaptive learning rate scheduler which is described in Appendix 8.1.
@@
The left panel of Figure 18 shows that the mean forecast time $\mathbb{E}[\tau_{f}]$ and the logarithm of the loss function $\log(\mathcal{L})$ are linearly related. This is a direct manifestation of the exponential sensitivity in chaotic dynamical systems: in each gradient descent step the loss experiences small changes leading to small changes in the learned weights and hence in the resulting surrogate model. These small changes in the chaotic surrogate model lead to an exponential divergence of nearby trajectories. This causes an exponential in time loss of predictability, characterized here by the mean of the forecast time (5).
@@
Figure 18. Left and right panels show the relationship between $\mathbb{E}[\tau_{f}]$ and the logarithm of the loss function $\mathcal{L}$ for a single-layer feedforward network and for a random feature map with only good internal parameters, i.e. $p_{g}=1$, respectively. Each dot in the left panel corresponds to a gradient descent step. Each dot in the right panel corresponds to one realization of a random feature map. The expectation is computed over $500$ validation trajectories. Each descent step and each realization use the same training and testing data. The black line in the left panel represents the best-fit line with slope $-0.79$. In the right panel the red crosses denote the conditional mean $\mathbb{E}[\tau_{f}|\log(\mathcal{L})]$ and the black line represents the best-fit line with slope $-1.01$. We use a feature dimension of $D_{r}=300$, training data length $N=20,000$ and regularization parameter $\beta=4\times 10^{-5}$
@@
The same sensitive dependency on small changes of the surrogate model, quantified by small changes of the loss function, is also present in random feature maps. The right panel of Figure 18 shows the mean forecast time $\mathbb{E}[\tau_{f}]$ as a function of the logarithm of the loss function for random feature maps. Each dot represents one realization of a random feature map with feature dimension $D_{r}=300$, trained on the same data as the single-layer feedforward network. The mean forecast time $\mathbb{E}[\tau_{f}]$ is computed using the same $500$ validation trajectories as the network. Averaged over bins of the logarithm of the loss function, the mean forecast shows the same linear relationship with respect to the logarithm of the loss function (red crosses in Figure 18). The slopes of the best-fit lines in Figure 18 show that the forecasting skill of the random feature map improves slightly faster with decreasing loss when compared to the network.
@@
The discrepancy between the neural network having worse forecasting skill compared to random feature maps despite achieving smaller loss can be explained as follows. Minimizing the loss function $\mathcal{L}$ aims at learning the single-step surrogate map (6). High forecasting skill, however, requires multiple applications of the single-step surrogate model which is not explicitly accounted for in the loss function (10). In Section 3.1 we established that the main controlling factor for achieving high forecasting skill is the number of good features. In random feature maps we can control and maximize this number simply by sampling good parameters according to our hit-and-run algorithms 1 and 2, respectively. On the other hand, the training of the single-layer feedforward network is only designed to minimize the loss but not to unable to maximize the number of good features to $N_{g}=D_{r}=300$ in our case. Figure 19 shows the number of different types of rows produced during the training instance of Figure 16. We see that only a single good row was produced in $(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in})$ during the early steps of the optimization, and this good row was then quickly destroyed during the training process. We checked that even when the network is initialized with only good internal parameters, i.e. $p_{g}=1$, training eventually leads to a complete absence of good internal parameters with $p_{g}=0$ for reasonable learning rates. To understand the absence of good rows in the trained network, note that for any random feature map $\mathbf{\Theta}=(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in},\mathbf{W})$ essentially lies on the graph of a continuous function due to the intricate relationship between the internal and outer weights dictated by (11). So the set of all possible $\mathbf{\Theta}$ for random feature maps has zero Lebesgue measure in $\mathbb{R}^{D_{r}\times(2D+1)}$. It is therefore highly unlikely that gradient descent finds the lower-dimensional subset of the random feature map weights in its search space which is the full $\mathbb{R}^{D_{r}\times(2D+1)}$. It would be interesting to see if the network generates good features if the loss function is augmented by a penalty term promoting good features. In any case, random feature maps are significantly cheaper to train.
@@
Figure 19. Evolution of the number (normalized by $D_{r}$) of learned good, linear and saturated rows $(\mathbf{w}_{i}^{\rm in},b_{i}^{\rm in})$ in the internal parameters during a single training episode of a single-layer feedforward network. The same neural network is used as in Figure 16. Results are shown every $10^{4}$ gradient descent steps.

### v2

@@ 5.1. Effect of the quality of internal weights on the forecast time $\tau_{f}$
In this section we investigate how the forecasting skill of a random feature surrogate model (3) improves with increasing number of good rows $N_{g}$. We would like to have internal parameters resulting in large mean forecast times $\tau_{f}$ with relatively small standard deviations. For chaotic dynamical systems we expect a residual variance of the forecast time due to the sensitivity to small changes in the model: small changes in the internal parameters may cause the surrogate models to deviate from each other after some time.
@@
Figure 9 shows the dependence of the mean forecast time $\mathbb{E}[\tau_{f}]$ and the associated coefficient of variation $\sigma[\tau_{f}]/\mathbb{E}[\tau_{f}]$ on $p_{g}$ for various values of the feature dimensions $D_{r}$ and training data lengths $N$. It is clearly seen that increasing the number of good rows increases the mean forecast time and decreases the coefficient of variation, as desired. As expected, for fixed feature dimension $D_{r}$ increasing the training data length $N$ is beneficial. A training data length of $N=78$ is too small to provide any meaningful forecasting skill with mean forecast times below one Lyapunov time for any value of $p_{g}$ and significantly larger variance. Similarly, for fixed training data length $N$, increasing the feature dimension $D_{r}$ is beneficial. The observation that, for fixed data length $N$, the mean forecast time $\mathbb{E}[\tau_{f}]$ saturates upon increasing $p_{g}$ once a sufficiently large number of good features $N_{g}=p_{g}D_{r}$ are present, suggests that the distribution of the forecast time $\tau_{f}$ converges reflecting a residual uncertainty of the chaotic surrogate model. This is confirmed in Figure 10 where we see convergence of the empirical histograms of the forecast time for increasing values of $D_{r}$ in the case when $p_{g}=1$.
@@
We briefly discuss the effect of the regularization parameter $\beta$ on the forecasting skill. We show in Figure 13 the mean forecast time $\mathbb{E}[\tau_{f}]$ and coefficient of variation $\sigma[\tau_{f}]/\mathbb{E}[\tau_{f}]$ as a function of $p_{g}$ for a range of regularization parameters $\beta\in[2^{-25},2^{-13}]$. For fixed feature dimension $D_{r}=300$, we see that $\beta=2^{-19}$ is optimal within this range in terms of the mean forecast time (left panel) and the coefficient of variation (right panel) once sufficiently many good features are present with $p_{g}>0.33$. Note that we had previously employed $\beta=4\times 10^{-5}\approx 2^{-14.6}$. The large difference in performance for different choices of the regularization parameter $\beta$ makes clear that, if optimizing for performance, the regularization parameter has to be optimized. This could be achieved by line-search [37] or by Bayesian optimization [26]. Here we have refrained from fine-tuning the regularization parameter as our focus is the sampling algorithm and the effect of different types of features on the forecast skill.
@@ 5.2. Effect of the quality of internal weights on the outer weights $\mathbf{W}$
We begin by recording the Frobenius norm $\mathbf{\|W\|}={\rm{Tr}}({\mathbf{W}}^{\top}{\mathbf{W}})$ of the learned outer weights as a function of the fraction of good features $p_{g}$ (bad features are again roughly equally distributed between linear and saturated features). Figure 14 shows the mean of $\mathbf{\|W\|}$ as a function of $p_{g}$ on a $\log$-$\log$ scale for the simulations used in Figure 9. It is seen that $\|\mathbf{W}\|$ is a decreasing function of the number of good features. The solution of the linear regression problem $\mathbf{W}$ minimizes the loss function (10). Once there are sufficiently many good features, the training data can be sufficiently well fit, decreasing the first term of the loss function. Further increasing the number of good features then allows for a decrease of the regularizing term of the loss function, leading to a decrease of $\|\mathbf{W}\|$. Assuming that the true one-step map $\Psi_{\Delta t}$ in (2) lies in the domain of the random feature map (3) with infinitely many features, the first term of the loss function should scale with the usual Monte-Carlo estimate scaling of $\mathcal{O}(1/D_{r})$, suggesting a scaling of the regularization term $\|\mathbf{W}\|\sim 1/\sqrt{D_{r}}$. In the right panel of Figure 14 we show that the mean of $\|\mathbf{W}\|$ roughly scales as $\|\mathbf{W}\|\sim 1/D_{r}^{0.54}$ when all the internal weights correspond to good features with $p_{g}=1$, suggesting that the true one-step map can be well approximated by random features with a $\tanh$-activation function. We remark that the Monte-Carlo scaling is valid for $D_{r}>256$ only, i.e. provided sufficiently many good features are present.
@@
Figure 14. Left: The mean of the Frobenius norm of the outer weights, $\|\mathbf{W}\|$, as a function of $p_{g}$ on a log-log scale. Right: The mean of the Frobenius norm of the outer weights, $\|\mathbf{W}\|$, as a function of the feature dimension $D_{r}$ for $p_{g}=1$. The line of best fit with approximate slope $-0.54$ is also shown on the right. The data are from the same experiments as shown in Figure 9.
@@
We now investigate how the decrease in the outer weights $\mathbf{W}$ is distributed over the various features. We will see that the outer weights are learned to suppress the bad features provided there are sufficiently many good features allowing for a reduction of the loss function. Let us denote the $i$-th column of $\mathbf{W}$ by $\mathbf{W}^{i}$. The columns $\mathbf{W}^{i}$ are the weights attributed to the features produced by the $i$-th row of the internal weights $(\mathbf{w}^{\rm in}_{i},b^{\rm in}_{i})$. We expect the outer weights corresponding to good rows to be significantly larger than those corresponding to bad rows.
@@
To study the suppression of bad features by columns of $\mathbf{W}$ which have small norm, we design two sets of numerical experiments: one in which bad features are entirely comprised of linear features and one in which bad features are entirely comprised of saturated features.
@@
In the first set we initialize a random feature map with $D_{r}=300$ features consisting of only bad linear features. We then successively replace one linear feature by a good feature, i.e. replacing one inner linear weight row $(\mathbf{w}_{i}^{\rm in},b_{i}^{\rm in})$ by a good row. At each step we record the corresponding linear regression solution $\mathbf{W}$. Figure 15 shows the normalized supremum norm of columns of $\mathbf{W}$ after $N_{g}=10$, $N_{g}=50$ and $N_{g}=150$ bad linear features have been replaced by good features. The red dots signify columns which do not contain any entry with absolute value larger than $1$. It is clearly seen that linear features are suppressed by the columns of $\mathbf{W}$. Note that not all linear features are entirely suppressed.
@@
In the second experiment, we follow the same procedure as before except we start with only saturated random features. In Figure 16 it is seen that saturated features are suppressed even stronger by the outer weights than linear features. In contrast to linear features, saturated features are effectively fully suppressed once the number of good features exceeds $N_{g}=50$.
@@
Figure 15. Normalized supremum norm of the columns of $\mathbf{W}$ for different numbers of good features with $N_{g}=10$, $N_{g}=50$ and $N_{g}=150$ and otherwise exclusively linear bad features. The $x$-axis represents column indices. The good and linear columns are indicated in blue and orange, respectively. The red dots signify columns with supremum norm less than $1$. The overall feature dimension is $D_{r}=300$ and the outer weights were obtained from training data of length $N=20,000$.
@@
Figure 16. Normalized supremum norm of the columns of $\mathbf{W}$ for different numbers of good features with $N_{g}=10$, $N_{g}=50$ and $N_{g}=150$ and otherwise exclusively saturated bad features. The $x$-axis represents column indices. The good and saturated columns are indicated in blue and green, respectively. The red dots signify columns with supremum norm less than $1$. The overall feature dimension is $D_{r}=300$ and the outer weights were obtained from training data of length $N=20,000$.
@@ 6. Comparison with a single-layer feedforward network trained with gradient descent
$\mathbf{{\color[rgb]{0,0,0}\Theta^{*}}}=\underset{\Theta}{\operatorname*{arg\,min}}\;\mathcal{L}({\color[rgb]{0,0,0}\Theta};\mathbf{U}),$ (30)
with $\mathbf{\Theta}=(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in},\mathbf{W})$ and the loss function $\mathcal{L}$ defined in (10). To solve the optimization problem (30) we employ gradient descent. To fairly compare with the results from the random feature model, we set the internal layer width to $D_{r}=300$, the regularization parameter to $\beta=4\times 10^{-5}$, and the training data length to $N=20,000$. To initialize the network weights we use the standard Glorot initialization [12]. We use an adaptive learning rate scheduler which is described in Appendix 9.1.
@@
The left panel of Figure 19 shows that the mean forecast time $\mathbb{E}[\tau_{f}]$ and the logarithm of the loss function $\log(\mathcal{L})$ are linearly related for the single-layer neural network. This is a direct manifestation of the exponential sensitivity in chaotic dynamical systems: in each gradient descent step the loss experiences small changes leading to small changes in the learned weights and hence in the resulting surrogate model. These small changes in the chaotic surrogate model lead to an exponential divergence of nearby trajectories. This causes an exponential in time loss of predictability, characterized here by the mean of the forecast time (5).
@@
Figure 19. Relationship between the mean forecast time $\mathbb{E}[\tau_{f}]$ and the logarithm of the loss function $\mathcal{L}$ for a single-layer feedforward network (left) and for a random feature map with only good internal parameters, i.e. $p_{g}=1$ (right). Each dot in the left panel corresponds to a gradient descent step. Each dot in the right panel corresponds to one realization of a random feature map. The expectation is computed over $500$ validation trajectories. Each descent step and each realization use the same training and testing data. The black line in the left panel represents the best-fit line. In the right panel the orange crosses denote the conditional mean $\mathbb{E}[\tau_{f}|\log(\mathcal{L})]$ and the black line represents the best-fit line. We use a feature dimension of $D_{r}=300$, training data length $N=20,000$ and regularization parameter $\beta=4\times 10^{-5}$.
@@
The same sensitive dependency on small changes of the surrogate model, quantified by small changes of the loss function, is also present in random feature maps. The right panel of Figure 19 shows the mean forecast time $\mathbb{E}[\tau_{f}]$ as a function of the logarithm of the loss function for random feature maps. Each dot represents one realization of a random feature map with feature dimension $D_{r}=300$, trained on the same data as the single-layer feedforward network. The mean forecast time $\mathbb{E}[\tau_{f}]$ is computed using the same $500$ validation trajectories as the network. Averaged over bins of the logarithm of the loss function, the mean forecast time shows the same linear relationship with the logarithm of the loss function (orange crosses in Figure 19). The slopes of the best-fit lines in Figure 19 show that the forecasting skill of the random feature map improves slightly faster with decreasing loss when compared to the single-layer network with an estimated slope of $-1.01$ for the random feature map and $-0.79$ for the neural network.
@@
The discrepancy between the neural network having worse forecasting skill compared to random feature maps despite achieving smaller loss can be explained as follows. Minimizing the loss function $\mathcal{L}$ aims at learning the single-step surrogate map (6). High forecasting skill, however, requires multiple applications of the single-step surrogate model which is not explicitly accounted for in the loss function (10). In Section 3.1 we established that the main controlling factor for achieving high forecasting skill is the number of good features. In random feature maps we can control and maximize this number simply by sampling good parameters according to our hit-and-run Algorithms 1 and 2, respectively. On the other hand, the training of the single-layer feedforward network is only designed to minimize the loss but not to maximize the number of good features to $N_{g}=D_{r}=300$ in our case. Figure 20 shows the number of different types of rows produced during the training instance of Figure 17. We see that only a single good row was produced in $(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in})$ during the early steps of the optimization, and this good row was then quickly destroyed during the training process. We checked that even when the network is initialized with only good internal parameters, i.e. $p_{g}=1$, training eventually leads to a complete absence of good internal parameters with $p_{g}=0$ for reasonable learning rates. To understand the absence of good rows in the trained network, note that for any random feature map $\mathbf{\Theta}=(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in},\mathbf{W})$ essentially lies on the graph of a continuous function due to the intricate relationship between the internal and outer weights dictated by (11). So the set of all possible $\mathbf{\Theta}$ for random feature maps has zero Lebesgue measure in $\mathbb{R}^{D_{r}\times(2D+1)}$. It is therefore highly unlikely that gradient descent finds the lower-dimensional subset of the random feature map weights in its search space which is the full $\mathbb{R}^{D_{r}\times(2D+1)}$. It would be interesting to see if the network generates good features if the loss function is augmented by a penalty term promoting good features. In any case, random feature maps are significantly cheaper to train.
@@
Figure 20. Evolution of the number (normalized by $D_{r}$) of learned good, linear and saturated rows $(\mathbf{w}_{i}^{\rm in},b_{i}^{\rm in})$ in the internal parameters during a single training episode of a single-layer feedforward network. The data shown here corresponds to the training instance shown in Figure 17 where we used $D_{r}=300$ and $N=20,000$. Results are shown every $10^{4}$ gradient descent steps.
7. Results for long-time statistical behaviour
We have so far focused on short-term integration and assessed the quality of a surrogate model by its ability to remain close to a reference trajectory. In certain applications such as climate prediction, however, it is more important to reliably recover the statistical properties of the dynamical system. Good short-term forecasting, i.e. high mean forecasting times, does not necessarily imply reliable reproduction of the statistics, and vice versa. In Figure 21 we compare the empirical histograms of the three variables of the Lorenz-63 system (4) with those of the corresponding random feature surrogate maps where we only used good features, i.e. $p_{g}=1$, and only used linear and saturated features, i.e. $p_{g}=0$ with $p_{l}=p_{s}=0.5$. The histograms are obtained from long simulations over $2,000$ time units, approximating the invariant measure. It is seen that the random feature map with only good features reproduces the invariant measure well. When only linear and saturated features are employed, the histogram of the Lorenz-63 system is less well reproduced. Remarkably, for $p_{g}=0$ the histogram is still reasonably well reproduced; in particular, the tail behaviour of the histograms is well reproduced.
We further show in Figure 22 a comparison between the histograms obtained from the random feature map surrogate model that used only good features with $p_{g}=1$, and the single-layer feedforward network. It is clearly seen that the neural network is not able to reproduce the long-time behaviour, with its empirical histogram being wildly different from that of the original Lorenz-63 system. However, the neural network also reproduces the tail behaviour of the histogram very well.
It is remarkable that including global information about the data, in the form of the constraints (13) which take into account information about the attractor, is sufficient to ensure that the trained surrogate model is able to recover the invariant measure. This makes random feature models with only good internal weights a very attractive network architecture for dynamical systems.
Figure 21. Empirical histograms of the marginals of the invariant measure for the Lorenz-63 system obtained from simulating the original dynamical system (4), and for random feature map surrogate model with only good features ($p_{g}=1$) and without any good features ($p_{g}=0$).
Figure 22. Empirical histograms of the marginals of the invariant measure for the Lorenz-63 system obtained from simulating the original dynamical system (4), for a random feature map surrogate model with $p_{g}=1$, and for a neural network (NN) surrogate model. The same neural network is used as in Figure 17.